OpenStudy (anonymous):

solve cos theta V3/2 for 0

OpenStudy (anonymous):

\(\cos(\theta)=\frac{\sqrt{3}}{2}\) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

look at the unit circle on the last page of the attached cheat sheet and find the two places where the first coordinate in the ordered pairs is \(\frac{\sqrt{3}}{2}\) then find the corresponding angle

OpenStudy (anonymous):

let me know what you get this is a very useful cheat sheet so i would save it, or if you lose it you can find it by googling "paul's notes" a great site for explanations and cheat sheets

OpenStudy (anonymous):

is it pi/6 and 5pi/6

OpenStudy (anonymous):

yes